### v1

@@ Abstract
This work develops user-friendly a posteriori error estimates of finite element methods, based on smoothers of linear iterative solvers. The proposed method employs simple smoothers, such as Jacobi or Gauss–Seidel iteration, on an auxiliary finer mesh to process the finite element residual for a posteriori error control. The implementation requires only a coarse-to-fine prolongation operator. For symmetric problems, we prove the reliability and efficiency of smoother-type error estimators under a saturation assumption. Numerical experiments for various PDEs demonstrate that the proposed smoother-type error estimators outperform residual-type estimators in accuracy and exhibit robustness with respect to parameters and polynomial degrees.
@@ 1 Introduction
Adaptive mesh refinement is a common routine in competitive numerical methods for resolving singularities and sharp gradients as well as saving computational cost in large-scale numerical simulations. In Adaptive Finite Element Methods (AFEMs), a posteriori error estimation is a key tool for equi-distributing errors and constructing nearly optimal locally refined meshes, see [52, 4, 53, 17] for basic ingredients of AFEMs and references therein. Classical a posteriori error estimates of AFEMs include residual, superconvergent recovery, dual weighted residual, equilibrated, hierarchical basis, functional a posteriori error estimators, see, e.g., [61, 5, 2, 15, 10, 24, 60, 49, 31, 44, 6, 40] for an incomplete list of references in this field.
@@
A posteriori error estimation of Finite Element Methods (FEMs) is also closely related to iterative linear solvers because both methods aim at developing computable approximations to the action of solution operators on some residuals. For instance, the idea of hierarchical basis solvers motivated the development of hierarchical basis error estimators in [8, 5]. Multigrid solvers combined with superconvergence are used for constructing a posteriori error estimates in [11, 12]. It has been shown in [41, 42] that additive operator preconditioning on infinite-dimensional spaces yields implicit and parameter-robust a posteriori error estimates for $H({\rm grad})$, $H({\rm curl})$ and $H({\rm div})$ problems. Conversely, a posteriori error estimation has been applied to design coarse spaces in multigrid methods, see [59, 36]
@@
We shall follow the research line in [41, 42] to derive a posteriori error estimates based on smoothers of iterative linear solvers. Smoothers are simple basic iterations such as Jacobi or Gauss–Seidel method on a fine mesh that are efficient for damping the high frequency component of the FEM error. We shall show that appropriate smoothers directly yield high-quality a posteriori error estimates of FEMs for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. In numerical experiments, smoother-based error estimators exhibit sharper effectivity ratio than residual-type a posteriori error estimates.
@@
The implementation of the smoother-type a posteriori error estimates on a mesh $\mathcal{T}_{h}$ is similar to a two-grid solver. In particular, we shall make use of a simple smoother $S$, e.g., a Jacobi, block Jacobi or Gauss–Seidel iteration on an auxiliary finer mesh $\mathcal{T}_{h/2}$. Then a smoother-type a posteriori error estimate is of the form $\|Sr\|$ or $\sqrt{\langle r,Sr\rangle}$, where $r$ is the finite element residual transferred to $\mathcal{T}_{h/2}$, $\|\bullet\|$ is the norm of interest and $\langle\bullet,\bullet\rangle$ the pairing between the primal and dual spaces. The reliability and efficiency of our a posteriori error estimates are confirmed under a saturation assumption.
@@
Smoother-type a posteriori error analysis in this paper offers a possibility that directly translates available iterative solvers of complex PDE systems into a posteriori error estimates inheriting many desirable properties of established solvers. For instance, complex multi-physics problems are modeled by saddle-point systems depending on a set of physical parameters. In the numerical study, a common first step is an discrete inf-sup stability analysis and a uniformly efficient block diagonal preconditioner. The proposed framework avoids redoing a posteriori error analysis for multi-physics systems and is able to rewrite their solvers as parameter-robust smoother-type a posteriori error estimates, see section 4.2 for an application in Biot’s consolidation model.
@@
In addition, the effectiveness of smoother-type a posteriori error estimates is robust with respect to the polynomial degree (referred to as $p$-robust) of finite elements. The $p$-robustness is an important property in $hp$-AFEMs and real-world simulations by very high order FEMs. In the literature, well-known $p$-robust a posteriori error estimates include equilibrated and superconvergent error estimators based on solving local problems on each vertex patch, see [19, 13]. Compared to classical $p$-robust estimators, pointwise smoother-type error estimators are easier to implement and do not necessitate solving auxiliary problems on local patches.
@@ 2 Abstract Framework
(2.1a) $a(u,v)$ $=f(v),\quad v\in V,$
(2.1b) $a(u_{h},v_{h})$ $=f(v_{h}),\quad v_{h}\in V_{h}.$
Throughout this paper, $V_{h}$ is a proper finite element space based on a partition $\mathcal{T}_{h}$ of the underlying domain $\Omega$ for (2.1a). The form $a$ incudes a bounded linear operator $A:V\rightarrow V^{*}$ defined as follows:
@@
Here $\langle\bullet,\bullet\rangle$ is the duality pairing between the primal and dual spaces. We say $A:V\rightarrow V^{*}$ is Symmetric and Positive-Definite (SPD) if $\langle Av,v\rangle\geq 0$ and $\langle Av,v\rangle=0\Leftrightarrow v=0$ for any $v\in V$. Clearly $A$ induces an inner product and a norm on $V$:
@@
We adopt the notation $C_{1}\lesssim C_{2}$ (resp. $C_{1}$ $\lesssim C_{2}$) provided $C_{1}\leq C_{3}C_{2}$ with $C_{3}$ being a generic constant independent of quantities of interest. We say $C_{1}\eqsim C_{2}$ provided $C_{1}\lesssim C_{2}$, $C_{2}\lesssim C_{1}$. This notation generalizes to comparison between SPD operators in an obvious way. It has been observed in [41] that additive two-level preconditioning yields a reliable and efficient error estimator for SPD problems. For completeness, we include a brief proof.
@@
$c_{0}\langle R,B_{a}R\rangle\leq\langle R,A^{-1}R\rangle\leq c_{1}\langle R,B_{a}R\rangle,\quad\forall R\in V^{*},$
@@
(2.2) $\sqrt{c}_{0}\langle r,S_{a}r\rangle^{\frac{1}{2}}\leq\|u-u_{h}\|_{A}\leq\sqrt{c}_{1}\langle r,S_{a}r\rangle^{\frac{1}{2}}.$
Proof 2.2.
@@
Noticing $\langle I_{h}^{*}r,v_{h}\rangle=\langle r,v_{h}\rangle=0~\forall v_{h}\in V_{h}$, we also obtain $B_{a}r=S_{a}r$. Then (2.2) follows from the spectral equivalence assumption $A^{-1}\eqsim B_{a}$.
If $V=V_{h}+\sum_{k=1}^{N}V_{k}$ is a stable decomposition, then the operator $B_{a}$ is a classical additive Schwarz preconditioner for $A$. In the literature, $S_{a}$ is referred to as an additive smoother for damping the high frequency component of the iterative error. Theorem 2.1 implies that the smoother $S_{a}$ yields a reliable and efficient a posteriori error estimate of finite element method. When The inverse $A_{k}^{-1}$ in $S_{a}$ often amounts to solving a local problem on $\Omega_{k}$.
We remark that $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ is applicable to SPD problems as well as symmetric but indefinite problems, see section 4. However, $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ could not be applied to non-symmetric problems because $\langle r,S_{a}r\rangle$ is not necessarily nonnegative in such cases. Our new observation is that a contractive iterative solver for $Au=f$
@@
Let $B:V^{*}\rightarrow V$ be a contractive iterator, i.e.,
$\|I-BA\|_{M}<1$
with respect to a norm $\|\cdot\|_{M}$ on $V$. Then with $r=f-Au_{h}$ we have
@@
(2.5) $\|BA\|_{M}\leq 1+\|I-BA\|_{M}<2.$
Combining (2.3) with (2.4) and (2.5) completes the proof.
@@
Let $S_{m}:V^{*}\rightarrow V$ be given in Algorithm 2.1. Assume that $V=V_{h}+\sum_{k=1}^{N}V_{k}$. Then the two-level iteration
(2.6a) $\eta$ $=u_{n}+A_{h}^{-1}I_{h}^{*}(f-Au_{n}),$
(2.6b) $u_{n+1}$ $=\eta+S_{m}(f-A\eta),\quad n=0,1,2,\ldots$
@@
Then contraction property of the algorithm (2.6) translates into $\|I-B_{m}A\|_{M}<1$. Due to $V_{h}\subset V$ and the definition of the Galerkin discretization (2.1b), we have $I_{h}^{*}r=0$ and thus $B_{m}r=S_{m}r.$ Using this identity and Theorem 2.3 completes the proof.
@@
(2.7) $-\Delta u$ $=f\quad\text{ in }\Omega,$
@@
(2.8a) $a(u,v)=(\nabla u,\nabla v)$ $=(f,v),\quad v\in H_{0}^{1}(\Omega),$
(2.8b) $(\nabla u_{h},\nabla v_{h})$ $=(f,v_{h}),\quad v_{h}\in V_{h},$
where $V_{h}\subset H_{0}^{1}(\Omega)$ is a finite element space of continuous and piecewise $\mathcal{P}_{1}$ polynomials. Let $\{a_{h,k}\}_{1\leq k\leq N_{h}}$ be the set of grid vertices in $\mathcal{T}_{h}$ and $\phi_{k}\in V_{h}$ the hat function associated to $a_{h,k}$. Then $\Omega=\cup_{k=1}^{N_{h}}\Omega_{h,k}$ with $\Omega_{h,k}={\rm supp}(\phi_{k})$ is an overlapping domain decomposition and $H_{0}^{1}(\Omega)=V_{h}+\sum_{k=1}^{N_{h}}H_{0}^{1}(\Omega_{h,k})$ is a stable space decomposition. In this case, the estimator $S_{a}$ in Theorem 2.1 leads to
(2.9) $|u-u_{h}|_{H^{1}(\Omega)}\eqsim\langle r,S_{a}r\rangle^{\frac{1}{2}}=\Big(\sum_{k=1}^{N_{h}}|\eta_{k}|_{H^{1}(\Omega_{h,k})}^{2}\Big)^{\frac{1}{2}},$
@@
(2.10) $|u-u_{h}|_{H^{1}(\Omega)}\eqsim|S_{m}r|_{H^{1}(\Omega)}.$

### v2

@@ Abstract
This work develops user-friendly a posteriori error estimates of finite element methods, based on smoothers of linear iterative solvers. The proposed method employs simple smoothers, such as Jacobi or Gauss–Seidel iteration, on an auxiliary finer mesh to process the finite element residual for a posteriori error control. The implementation has linear complexity and requires only a coarse-to-fine prolongation operator. For symmetric problems, we prove the reliability and efficiency of smoother-type error estimators under a saturation assumption. Numerical experiments for various PDEs demonstrate that the proposed smoother-type error estimators outperform residual-type estimators in accuracy and exhibit robustness with respect to parameters and polynomial degrees.
@@
journal: Computer Methods in Applied Mechanics and Engineering
[…]
[1]organization=School of Mathematical Sciences, Zhejiang University, addressline=866 Yuhangtang Road, city=Hangzhou, postcode=310058, state=Zhejiang, country=People’s Republic of China
22footnotetext: Funding:  This work was supported by the National Key R&D Program of China under grant 2024YFA1012600 and the National Natural Science Foundation of China under grant 12471346.
@@ 1 Introduction
Adaptive mesh refinement is a common routine in competitive numerical methods for resolving singularities and sharp gradients as well as saving computational cost in large-scale numerical simulations. In Adaptive Finite Element Methods (AFEMs), a posteriori error estimation is a key tool for equi-distributing errors and constructing nearly optimal locally refined meshes, see [61, 5, 62, 19] for basic ingredients of AFEMs and references therein. Classical a posteriori error estimates of AFEMs include residual, superconvergent recovery, dual weighted residual, equilibrated, hierarchical basis, functional a posteriori error estimators, see, e.g., [70, 6, 3, 17, 11, 26, 69, 2, 58, 36, 37, 50, 54, 7, 49] for an incomplete list of references in this field.
@@
A posteriori error estimation of Finite Element Methods (FEMs) is also closely related to iterative linear solvers because both methods aim at developing computable approximations to the action of solution operators on some residuals. For instance, the idea of hierarchical basis solvers motivated the development of hierarchical basis error estimators in [9, 6]. Multigrid solvers combined with superconvergence are used for constructing a posteriori error estimates in [12, 13]. It has been shown in [51, 52] that additive operator preconditioning on infinite-dimensional spaces yields implicit and parameter-robust a posteriori error estimates for $H({\rm grad})$, $H({\rm curl})$ and $H({\rm div})$ problems. Conversely, a posteriori error estimation has been applied to design coarse spaces in multigrid methods and to control iterative algebraic errors in domain decomposition methods, see [68, 44, 15].
@@
We shall follow the research line in [51, 52] to derive a posteriori error estimates based on smoothers of iterative linear solvers. Smoothers are simple basic iterations such as Jacobi or Gauss–Seidel method on a fine mesh and are efficient for damping the high frequency component of the FEM error. We shall show that appropriate smoothers directly yield high-quality a posteriori error estimates of FEMs for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. In numerical experiments, smoother-based error estimators exhibit sharper effectivity ratio than residual-type a posteriori error estimates. In contrast to the additive preconditioning-type implicit error estimators derived in [51, 52], which involve incomputable local problems, the proposed smoother-type a posteriori error estimators are fully computable. Moreover, this work introduces a new family of error estimators inspired by contractive iteration methods, which are particularly well-suited for handling non-symmetric problems.
@@
The implementation of the smoother-type a posteriori error estimates on a mesh $\mathcal{T}_{h}$ is similar to a two-grid solver. In particular, we shall make use of a simple smoother $S$, e.g., a Jacobi, block Jacobi or Gauss–Seidel iteration on an auxiliary finer mesh $\mathcal{T}_{h/2}$. Then a smoother-type a posteriori error estimate is of the form $\|Sr\|$ or $\sqrt{\langle r,Sr\rangle}$, where $r$ is the finite element residual transferred to $\mathcal{T}_{h/2}$, $\|\bullet\|$ is the norm of interest and $\langle\bullet,\bullet\rangle$ the pairing between the primal and dual spaces. The reliability and efficiency of our a posteriori error estimates are confirmed under a saturation assumption. Using the same assumption, adaptive finite and boundary element methods driven by the difference $\|u_{h}-u_{h/2}\|$ of finite element solution on nested meshes have been analyzed in [39, 38], while computing finer mesh solution $u_{h/2}$ is much more expensive than the smoother action $Sr$.
@@
Smoother-type a posteriori error analysis in this paper offers a possibility that directly translates available iterative solvers of complex PDE systems into a posteriori error estimates inheriting many desirable properties of established solvers. For instance, complex multi-physics problems are modeled by saddle-point systems depending on a set of physical parameters (e.g., Lamé constants and permeability). In the numerical study, a common first step is a discrete inf-sup stability analysis and a uniformly efficient block diagonal preconditioner. The proposed framework avoids redoing a posteriori error analysis for multi-physics systems and is able to rewrite their solvers as parameter-robust smoother-type a posteriori error estimates, see section 4.2 for an application in Biot’s consolidation model.
@@
In addition, numerical experiments demonstrate that the effectiveness of smoother-type a posteriori error estimates is robust with respect to the polynomial degree (referred to as $p$-robust) of finite elements. The $p$-robustness is an important property in $hp$-AFEMs and real-world simulations by very high order FEMs (cf. [61, 30, 31]). In the literature, well-known $p$-robust a posteriori error estimates include equilibrated and superconvergent error estimators based on solving local problems on each vertex patch, see [21, 14]. Compared to classical $p$-robust estimators, pointwise smoother-type error estimators are easier to implement and do not necessitate solving auxiliary problems on local patches.
@@ 2 Abstract Framework
In this section, we present two smoother-type a posteriori error estimates for an abstract positive-definite model posed in a Hilbert space that admits a stable subspace decomposition.
2.1 Abstract Error Estimators
@@
$a(u,v)$ $=f(v),\quad v\in V,$ (2.1a)
$a(u_{h},v_{h})$ $=f(v_{h}),\quad v_{h}\in V_{h}.$ (2.1b)
Throughout this paper, $V_{h}$ is a proper finite element space based on a partition $\mathcal{T}_{h}$ of the underlying domain $\Omega$ for (2.1a). The form $a$ includes a bounded linear operator $A:V\rightarrow V^{*}$ defined as follows:
@@
Here $\langle\bullet,\bullet\rangle$ is the duality pairing between the primal and dual spaces. We say $A:V\rightarrow V^{*}$ is Symmetric and Positive-Definite (SPD) if $\langle Av,v\rangle\geq 0$ and $\langle Av,v\rangle=0\Longleftrightarrow v=0$ for any $v\in V$. Clearly $A$ induces an inner product and a norm on $V$:
@@
We adopt the notation $C_{1}\lesssim C_{2}$ (resp. $C_{1}$ $\lesssim C_{2}$) provided $C_{1}\leq C_{3}C_{2}$ with $C_{3}$ being a generic constant independent of quantities of interest, e.g., the mesh size parameter $h$ and target functions. We say $C_{1}\eqsim C_{2}$ provided $C_{1}\lesssim C_{2}$, $C_{2}\lesssim C_{1}$. This notation generalizes to comparison between SPD operators in an obvious way. It has been observed in [51] that additive two-level preconditioning yields a reliable and efficient error estimator for SPD problems. For completeness, we include a brief proof.
@@
$c_{0}\langle R,B_{a}R\rangle\leq\langle R,A^{-1}R\rangle\leq c_{1}\langle R,B_{a}R\rangle,\quad\forall R\in V^{*},$ (2.2)
@@
$\sqrt{c}_{0}\langle r,S_{a}r\rangle^{\frac{1}{2}}\leq\|u-u_{h}\|_{A}\leq\sqrt{c}_{1}\langle r,S_{a}r\rangle^{\frac{1}{2}}.$ (2.3)
Proof.
@@
Noticing $\langle I_{h}^{*}r,v_{h}\rangle=\langle r,v_{h}\rangle=0~\forall v_{h}\in V_{h}$, we also obtain $B_{a}r=S_{a}r$. Then (2.3) follows from the spectral equivalence assumption $A^{-1}\eqsim B_{a}$. ∎
If $V=V_{h}+\sum_{k=1}^{N}V_{k}$ is a stable subspace decomposition, i.e., for any $v\in V$, there exist $v_{h}\in V_{h}$ and $v_{k}\in V_{k}$ with $k=1,\ldots,N$ such that
${v=v_{h}+\sum_{k=1}^{N}v_{k},}$
${\|v_{h}\|_{A}^{2}+\sum_{k=1}^{N}\|v_{k}\|_{A}^{2}\lesssim\|v\|_{A}^{2},}$
then the spectral equivalence (2.2) holds and $B_{a}$ is a classical additive Schwarz preconditioner for $A$. The spaces $H^{1}(\Omega)$, $H({\rm curl},\Omega)$, $H({\rm div},\Omega)$ and their finite element discrete variants under consideration all admit an appropriate stable decomposition for the analysis. In the literature, $S_{a}$ is referred to as an additive smoother for damping the high frequency component of the iterative error. Theorem 2.1 implies that the smoother $S_{a}$ yields a reliable and efficient a posteriori error estimate of finite element method. The inverse $A_{k}^{-1}$ in $S_{a}$ often amounts to solving a local problem on $\Omega_{k}$.
We remark that $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ is applicable to SPD problems as well as symmetric but indefinite systems posed in a Cartesian product space $V_{1}\times V_{2}\times\cdots$, where $S_{a}={\rm diag}(S_{V_{1}},S_{V_{2}},\ldots)$ is block diagonal with each $S_{V_{i}}$ being a SPD smoother on $V_{i}$, see Section 4. However, $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ could not be applied to strongly non-symmetric problems because $\langle r,S_{a}r\rangle$ is not necessarily nonnegative in such cases. Our new observation is that a contractive iterative solver for $Au=f$
@@
Let $\|\bullet\|_{M}$ be a norm on $V$. Let $B:V^{*}\rightarrow V$ be a contractive iterator with respect to the $M$-norm, i.e.,
$\|I-BA\|_{M}<1.$
Then with $r=f-Au_{h}\in V^{*}$, we have
@@
$\|BA\|_{M}\leq 1+\|I-BA\|_{M}<2.$ (2.6)
Combining (2.4) with (2.5) and (2.6) completes the proof. ∎
Remark 2.3.
In most cases, $\|\bullet\|_{M}$ is taken to be the underlying norm $\|\bullet\|_{A}$ induced by the SPD operator $A$. We note, however, that $\|\bullet\|_{M}$ practically need not coincide with the natural norm $\|\bullet\|_{A}$. For instance, choosing $\|\bullet\|_{M}=\|\bullet\|_{L^{2}(\Omega)}$ yields a posteriori error estimates in the $L^{2}$-norm for problems posed in $V=H^{1}(\Omega)$ or $V=H({\rm curl},\Omega)$; see Sections 3.1.3 and 3.2.
@@
Let $\|\bullet\|_{M}$ be a norm on $V$ and $S_{m}:V^{*}\rightarrow V$ be given in Algorithm 2.1. Assume that $V=V_{h}+\sum_{k=1}^{N}V_{k}$. Then the two-level iteration
$\eta$ $=u_{n}+A_{h}^{-1}I_{h}^{*}(f-Au_{n}),$ (2.7a)
$u_{n+1}$ $=\eta+S_{m}(f-A\eta),\quad n=0,1,2,\ldots$ (2.7b)
@@
Then contraction property of the algorithm (2.7) translates into $\|I-B_{m}A\|_{M}<1$. Due to $V_{h}\subset V$ and the definition of the Galerkin discretization (2.1b), we have $I_{h}^{*}r=0$ and thus $B_{m}r=S_{m}r.$ Using this identity and Theorem 2.2 completes the proof. ∎
When AA is the SPD operator induced by the inner product of $H^{1}(\Omega)$, $H({\rm curl},\Omega)$, $H({\rm div},\Omega)$ or their finite element variants, the iteration (2.7) with $S_{m}$ in Algorithm 2.1 given by a stable decomposition is uniformly contractive under the norm $\|\bullet\|_{A}$, i.e., $\|I-B_{m}A\|_{A}<1$ is independent of the mesh size parameter $h$.
2.2 Application to the Poisson Equation
@@
$-\Delta u$ $=f\quad\text{ in }\Omega,$ (2.8)
@@
$a(u,v)=(\nabla u,\nabla v)$ $=(f,v),\quad v\in H_{0}^{1}(\Omega),$ (2.9a)
$(\nabla u_{h},\nabla v_{h})$ $=(f,v_{h}),\quad v_{h}\in V_{h},$ (2.9b)
where $V_{h}\subset H_{0}^{1}(\Omega)$ is a finite element space of continuous and piecewise $\mathcal{P}_{1}$ polynomials. Let $\{a_{h,k}\}_{1\leq k\leq N_{h}}$ be the set of grid vertices in $\mathcal{T}_{h}$ and $\phi_{h,k}\in V_{h}$ the hat function associated to $a_{h,k}$. Then $\Omega=\cup_{k=1}^{N_{h}}\Omega_{h,k}$ with $\Omega_{h,k}={\rm supp}(\phi_{h,k})$ is an overlapping domain decomposition and $H_{0}^{1}(\Omega)=V_{h}+\sum_{k=1}^{N_{h}}H_{0}^{1}(\Omega_{h,k})$ is a stable subspace decomposition. In this case, the estimator $S_{a}$ in Theorem 2.1 leads to
$|u-u_{h}|_{H^{1}(\Omega)}\eqsim\langle r,S_{a}r\rangle^{\frac{1}{2}}=\Big(\sum_{k=1}^{N_{h}}|\eta_{k}|_{H^{1}(\Omega_{h,k})}^{2}\Big)^{\frac{1}{2}},$ (2.10)
@@
$|u-u_{h}|_{H^{1}(\Omega)}\eqsim|S_{m}r|_{H^{1}(\Omega)}.$ (2.11)